# Can Large Language Models Predict the Outcome of Judicial Decisions?

### v1

@@ Abstract
Large Language Models (LLMs) have shown exceptional capabilities in Natural Language Processing (NLP) across diverse domains. However, their application in specialized tasks such as Legal Judgment Prediction (LJP) for low-resource languages like Arabic remains underexplored. In this work, we address this gap by developing an Arabic LJP dataset, collected and preprocessed from Saudi commercial court judgments. We benchmark state-of-the-art open-source LLMs, including LLaMA-3.2-3B and LLaMA-3.1-8B, under varying configurations such as zero-shot, one-shot, and fine-tuning using QLoRA. Additionally, we used a comprehensive evaluation framework combining quantitative metrics (BLEU and ROUGE) and qualitative assessments (Coherence, legal language, clarity). Our results demonstrate that fine-tuned smaller models achieve comparable performance to larger models in task-specific contexts while offering significant resource efficiency. Furthermore, we investigate the effects of prompt engineering and fine-tuning on model outputs, providing insights into performance variability and instruction sensitivity. By making the dataset, implementation code, and models publicly available, we establish a robust foundation for future research in Arabic legal NLP.
@@
LLMs have transformed NLP, achieving state-of-the-art (SOTA) performance across diverse tasks in multilingual and multitask settings [1]. Advanced LLMs like Generative Pre-trained Transformer 4 (GPT-4) and Gemini demonstrate remarkable reasoning, comprehension, and problem-solving capabilities, enabling applications in specialized domains such as law [2], medicine [3, 4], and education [5].
However, their closed-source nature raises concerns about data privacy, customization, and accessibility, especially for organizations seeking domain-specific solutions [6, 7]. Open-source models like Large Language Model Meta AI (LLaMA)-3 [8] and Philosophically Integrated-4 (Phi-4) [9] provide flexible alternatives, offering fine-tuning and in-context learning capabilities that often rival proprietary systems and exceed baseline results [10, 11].
A key application of LLMs is LJP, which predicts judicial outcomes based on factual case details [12]. Currently, legal experts rely on extensive training to identify relevant laws, determine charges, and issue judgments [13]. Legal decision-making is inherently complex, requiring analysis of vast case-specific information while ensuring consistency and fairness [14]. This process remains labor-intensive, time-consuming, and prone to bias [14].
@@
To address these challenges, this work develops a tailored solution for Arabic Legal Judgment Prediction. Specifically, we:
@@
Collect and preprocess Arabic legal case data from the Saudi commercial court111https://laws.moj.gov.sa/ar/, creating a domain-specific dataset.
@@
Craft 75 diverse Arabic instructions of varying lengths and styles to simulate user-centric prompting scenarios.
@@
Benchmark open-source LLMs, including LLaMA-3.2-3B-instruct and LLaMA-3.1-8B-instruct, under zero-shot, one-shot, and fine-tuning settings using QLoRA.
@@ II Related Work
Recent years have seen significant interest in LJP, particularly with the rise of Machine Learning (ML) and NLP techniques. Early approaches to LJP primarily relied on rule-based systems and manually crafted features. While these methods were foundational, they were constrained by scalability and efficiency challenges [18]. The introduction of transformer-based models, such as the Bidirectional Encoder Representations from Transformers (BERT) model [19], marked a turning point by enabling models to capture complex semantic and syntactic relationships in legal texts. For example, Imran et al. [20] demonstrated BERT’s effectiveness in classifying European Court of Human Rights cases, showcasing its capacity to process intricate legal documents. Building on the success of transformers, researchers turned to ML-based approaches that emphasized data-driven learning, further enhancing the ability of models to handle unstructured legal texts. Similarly, ALJP [21] focused on Arabic legal texts, addressing challenges like morphological complexity and demonstrating that deep learning approaches outperformed traditional classifiers in this domain.
Aletras et al. [22] conducted one of the first studies on predicting case outcomes for the European Court of Human Rights (ECHR). Using a Support Vector Machine (SVM) [23] trained on N-grams and topic features, they achieved 79% accuracy in determining whether a human rights article had been violated. This pioneering work demonstrated the potential of using textual data for judicial decision-making and inspired further research into automated legal outcome prediction. However, the reliance on manual features in such approaches highlighted the need for more robust and scalable methods.
II-A Transformer-Based Models in Legal NLP
Transformer-based models revolutionized NLP by enabling the modeling of complex semantic and syntactic relationships. The BERT model [19] and its variants have significantly advanced legal text analysis by capturing semantic richness.
[20] demonstrated BERT’s effectiveness in classifying European Court of Human Rights cases, highlighting its ability to handle intricate legal documents. Similarly, ALJP: An Arabic Legal Judgment Prediction tackled challenges in Arabic legal texts, such as morphological richness and syntax complexity, showing that deep learning approaches, including transformers, outperformed traditional classifiers like SVM and Random Forest [21]. AraBERT, a transformer model fine-tuned for Arabic, has also delivered strong performance across Arabic NLP tasks, emphasizing its value for domain-specific applications like LJP [24]. Collectively, these studies demonstrate how transformer models have driven significant advancements in legal NLP, particularly in case classification and legal document understanding.
II-B ML-Based Approaches to LJP
As ML methodologies progressed, researchers shifted their focus toward data-driven methods leveraging increasingly accessible law-related datasets. This shift represented a crucial evolution in LJP research, allowing models to learn directly from extensive legal document collections rather than relying on manually crafted features.
Cui et al. [25] explored the impact of NLP techniques on processing unstructured legal documents, facilitating tasks like legal text classification, summarization, and case outcome prediction. Additionally, specialized transformer architectures like LegalBERT [26] and GPT have proven particularly effective in handling the intricate structure of legal texts, enabling advancements in legal-specific tasks. However, most of these models focus on English-language datasets, with limited research on applying such techniques to Arabic’s unique linguistic challenges [27].
II-C Text-Generation Techniques for Legal NLP
Text-generation techniques have become transformative in legal NLP, supporting tasks like document summarization, automated contract drafting, and case argument generation. Large pre-trained language models like GPT-3 and GPT-4 have made generating coherent and contextually accurate legal text increasingly feasible. For instance, [28] evaluated GPT-3’s performance on the Uniform Bar Examination (UBE), a standardized test for legal professionals, showcasing its ability to comprehend legal reasoning and provide contextually relevant answers in multiple-choice sections. Similarly, transformer-based models like GPT have demonstrated their capacity to process and generate structured, long-form legal text, highlighting their potential in specialized applications such as drafting judgments.
II-D LLM-Based Performance for Low-Resource Languages
LLMs excel in high-resource languages like English but face challenges in low-resource languages, such as Arabic, due to pre-training on English-heavy datasets. Conneau et al. [29] demonstrated that multilingual models like XLM-R can generalize across languages, handling cross-lingual tasks with limited data. Similarly, Pires et al. [30] showed that multilingual BERT effectively transfers knowledge across languages, achieving strong results in low-resource tasks through cross-lingual embeddings.
@@ II-D LLM-Based Performance for Low-Resource Languages
Interestingly, English-written instructions often yield strong performance even for non-English tasks. For instance, [17] found that non-native English instructions surpassed native Arabic ones in certain Arabic NLP tasks, highlighting the adaptability of LLMs when leveraging high-quality, generalized instruction data in low-resource settings.
II-E Research Gap in Arabic Legal NLP
Despite advancements in legal NLP and the rise of transformer-based models, notable gaps persist in Arabic LJP. A major obstacle is the absence of publicly available Arabic legal datasets. Existing corpora are often fragmented, domain-specific, or insufficiently annotated, limiting the ability to train accurate models for Arabic LJP. The lack of high-quality, large-scale datasets has slowed the development of effective solutions for the Arabic legal domain.
Additionally, Arabic legal NLP lacks benchmarking frameworks akin to those available for high-resource languages like English. This shortfall hinders objective evaluation and comparison of models and prompting techniques. Establishing evaluation frameworks tailored to Arabic legal tasks, such as LJP, is essential to advance the field, ensure reproducibility, and foster further research.
To the best of our knowledge,This work introduces the first publicly available Arabic Legal Judgment Prediction dataset and model. While prior studies have explored Arabic NLP and some legal tasks, research focused on Arabic LJP remains scarce. By providing a curated dataset and a domain-specific model, this study addresses a critical gap and establishes a foundation for progress in Arabic legal NLP.
Figure 1 outlines the complete methodology, starting from dataset collection from the Saudi Ministry of Justice, through data preparation and instruction crafting, to experimental setups involving fine-tuning and evaluation. It highlights the systematic process for benchmarking models like LLaMA-3.2-3B and LLaMA-3.1-8B on Arabic Legal Judgment Prediction.
@@ III-A Dataset Collection and Preparation
In this research, we constructed a domain-specific dataset for Arabic LJP by collecting court judgments from the publicly accessible Saudi Ministry of Justice Judgment Publication Platform222https://laws.moj.gov.sa/ar/. This platform hosts a substantial repository of legal cases, particularly within the domain of commercial law. Fig. 2 presents a sample data point.
@@
The dataset is stored as JSON files, with each data point containing several fields: a unique identifier (id), a reference to the original legal case (original_id), the concatenated facts and reasons from the case (input), the court’s final judgment (output), the dataset name (e.g., LJP), the task type (Legal Judgment Prediction), the language (Ar), a detailed task instruction (Instruction), and a numerical vector representation of the input (embedding).
The embeddings were generated using a BERT-based sentence embedding model from Sentence Transformers [31], specifically the paraphrase-multilingual-MiniLM-L12-v2 model333https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2. This model supports multiple languages, including Arabic, and enables efficient semantic retrieval of relevant legal judgments.
This following table illustrates how the model processes legal instructions and facts to predict judicial outcomes. It highlights the importance of clear prompts and structured data in enhancing LLM performance for tasks like Arabic LJP.
III-B QLoRA Fine-Tuning
To fine-tune our model efficiently, we adopt QLoRA (Quantized Low-Rank Adaptation) [32], a parameter-efficient fine-tuning technique that enables the adaptation of large pre-trained models without a significant increase in computational cost. QLoRA combines two key principles: quantization and low-rank adaptation. It quantizes the model’s parameters into 4-bit precision, significantly reducing memory usage while preserving performance. This allows fine-tuning on resource-limited hardware, making it viable for large-scale applications.
The core mechanism of QLoRA builds upon LoRA (Low-Rank Adaptation) [33], which factorizes the weight matrix into two low-rank matrices, as follows:
where , , and is the rank of the factorization, with and . Here, represents the low-rank modification to , allowing parameter-efficient updates.
During training, the original weights remain frozen and do not receive gradient updates, ensuring that the pre-trained knowledge is preserved. Only the low-rank matrices and are trainable. To control the contribution of the LoRA modifications, is scaled by a factor , where is a scaling hyperparameter. Higher values of amplify the influence of the low-rank updates, enabling more expressive adaptations.
The experimental setup uses a batch size of 2, gradient accumulation steps of 2, a maximum sequence length of 2048, LoRA rank , and a scaling factor . The target modules include q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, and down_proj, with no dropout applied. Training is conducted over one epoch using a learning rate of and the AdamW optimizer [34].
@@ III-C Evaluation Metrics
For quantitative evaluation, we employed ROUGE-F [35] and BLEU [36], which are standard metrics for text generation tasks:
@@ III-E Model Inference
The models were deployed for inference using vLLM [39] on two NVIDIA A16 GPUs. The maximum sequence length was set to 2048 tokens. To ensure reproducibility and deterministic outputs, the temperature parameter was fixed at zero.
@@ IV Experiments and Results
IV-A Subjective Metrics
The LLaMA-3.1-8B-Instruct achieves the highest scores in coherence (5.49), legal language (6.69), and faithfulness (5.99), demonstrating the advantages of increased parameter size. The LLaMA-3.2-3B-Instruct-Finetuned performs competitively, excelling in clarity (5.94) and consistency (6.16), underscoring the benefits of task-specific fine-tuning. The one-shot model improves over the base model across all metrics but lags behind the fine-tuned and larger models.
IV-B Objective Metrics
For ROUGE-1, ROUGE-2, and BLEU, the LLaMA-3.2-3B-Instruct-Finetuned achieves the highest scores (0.34, 0.25, 0.14), demonstrating strong alignment with reference outputs. The one-shot model performs moderately well, while the LLaMA-3.1-8B-Instruct underperforms, indicating that fine-tuning provides more significant gains than scale alone for these metrics.
IV-C Comparative Insights
The LLaMA-3.1-8B-Instruct emerges as the strongest general-purpose model, but the LLaMA-3.2-3B-Instruct-Finetuned achieves comparable or superior performance in task-specific contexts with reduced computational cost. While prompting improves performance, it remains less effective than fine-tuning for domain-specific tasks.
Fine-tuning smaller models can surpass larger models in specialized tasks, offering a resource-efficient alternative.
| Metric | LLaMA-3.2-3B | LLaMA-3.1-8B | LLaMA-3.2-3B-1S | LLaMA-3.2-3B-FT |
| Coherence | 2.69 | 5.49 | 4.52 | 5.31 |
| Brevity | 1.99 | 4.30 | 3.76 | 4.98 |
| Legal Language | 3.66 | 6.69 | 5.18 | 5.87 |
| Faithfulness | 3.00 | 5.99 | 4.00 | 5.24 |
| Clarity | 2.90 | 5.79 | 4.99 | 5.94 |
| Consistency | 3.04 | 5.93 | 5.14 | 6.16 |
| Avg. Qualitative Score | 3.01 | 5.89 | 4.66 | 5.68 |
| ROUGE-1 | 0.08 | 0.12 | 0.29 | 0.34 |
| ROUGE-2 | 0.02 | 0.04 | 0.19 | 0.25 |
| BLEU | 0.01 | 0.02 | 0.11 | 0.14 |
@@ V Discussion
This section examines the performance of LLM configurations in LJP, focusing on the effects of prompt engineering, fine-tuning, and in-context learning. We analyze model behaviors, highlight strengths and limitations, and discuss trade-offs in adapting LLMs for this specialized task.
V-A Prompt Engineering Effect for LJP
Statistical analyses using one-way ANOVA and the Kruskal-Wallis test were conducted to assess how prompt instructions affected model performance in LJP. The goal was to determine if significant variability existed across different instructions.
The Llama-3.2-3b-1shot model displayed no significant differences across instructions (ANOVA: , ; Kruskal-Wallis: , ), indicating consistent responses regardless of prompt variations. In contrast, Llama-3.2-3b-finetuned showed significant differences in ANOVA (, ) but not in the Kruskal-Wallis test (, ), suggesting fine-tuning increases sensitivity to instruction variability.
The Llama-3.2-3b-instruct model showed marginal evidence of inconsistency (ANOVA: , ; Kruskal-Wallis: , ). Although not statistically significant, this highlights the subtle effects prompt variability may have on instruction-tuned models [40]. On the other hand, the Llama-3.1-8b-instruct model demonstrated robust consistency (ANOVA: , ; Kruskal-Wallis: , ), likely due to its larger size and diverse instruction pre-training, aligning with scaling laws for language models [41].
These findings reveal key patterns in prompt engineering effects. Smaller models, such as Llama-3.2-3b-1shot, perform consistently but exhibit lower overall performance. Fine-tuning smaller models, as with Llama-3.2-3b-finetuned, significantly improves performance, comparable to larger models, but increases sensitivity to instruction structure. Larger models, like Llama-3.1-8b-instruct, perform better overall and show reduced dependence on prompt quality.
Effective prompts that are specific, structured, and contextually aligned with legal reasoning consistently enhance model performance, enabling accurate and coherent outputs. Conversely, vague or overly broad prompts lead to less relevant and disjointed responses. These results highlight the critical role of prompt clarity and structure in optimizing LJP tasks.
V-B Effect of Fine-tuning and In-Context Learning
Table I shows significant ROUGE-2 and BLEU score improvements for the LLaMA-3.2-3B model with 1-shot learning and fine-tuning compared to the base model (LLaMA-3.2-3B-Instruct). Specifically, 1-shot learning, as an in-context learning approach, achieves an 850% improvement in ROUGE-2 (0.02 to 0.19) and a 1000% improvement in BLEU (0.01 to 0.11). While enhancing response quality, style, and correctness, it falls short of fine-tuning. Fine-tuning delivers greater gains, with a 1150% increase in ROUGE-2 (0.02 to 0.25) and a 1300% increase in BLEU (0.01 to 0.14), enriching task-specific knowledge and generating more accurate, relevant responses. These results highlight fine-tuning’s effectiveness in task adaptation and the lighter, albeit less impactful, benefits of in-context learning.
V-C Limitations
While this study provides valuable insights into the effects of prompt engineering and fine-tuning on LLM performance in LJP, several limitations should be noted:
Reliance on LLM-Based Evaluation:
We relied on an LLM to score model outputs against the ground truth, However, the evaluation’s accuracy and reliability were not systematically validated, risking potential biases or misinterpretations.
Limited Model Size:
Computational constraints limited us to smaller models, thoughLarger models are known to yield better performance across multiple NLP tasks. Exploring the impact of scaling remains a future direction.
Restricted Number of Shots:
We used a 1-shot setting for in-context learning to balance performance and efficiency. Increasing the shots could improve performance but was not feasible due to resource constraints.
QLoRA Fine-Tuning:
Fine-tuning was performed on a 4-bit quantized model using QLoRA, which is efficient but may under-perform compared to full-model fine-tuning.
Addressing these limitations in future work will deepen understanding of LLM strengths and weaknesses in LJP.

### v3

@@ Abstract
Large Language Models (LLMs) have shown exceptional capabilities in Natural Language Processing (NLP) across diverse domains. However, their application in specialized tasks such as Legal Judgment Prediction (LJP) for low-resource languages like Arabic remains underexplored. In this work, we address this gap by developing an Arabic LJP dataset, collected and preprocessed from Saudi commercial court judgments. We benchmark state-of-the-art open-source LLMs, including LLaMA-3.2-3B and LLaMA-3.1-8B, under varying configurations such as zero-shot, one-shot, and fine-tuning using LoRA. Additionally, we employed a comprehensive evaluation framework that integrates both quantitative metrics (such as BLEU, ROUGE, and BERT) and qualitative assessments (including Coherence, Legal Language, Clarity, etc.) using an LLM. Our results demonstrate that fine-tuned smaller models achieve comparable performance to larger models in task-specific contexts while offering significant resource efficiency. Furthermore, we investigate the impact of fine-tuning the model on a diverse set of instructions, offering valuable insights into the development of a more human-centric and adaptable LLM. We have made the dataset, code, and models publicly available to provide a solid foundation for future research in Arabic legal NLP111https://github.com/MohamedBayan/Arabic-Legal-Judgment-Prediction.
@@
LLMs have transformed NLP, achieving state-of-the-art (SOTA) performance across diverse tasks in multilingual and multitask settings [1]. Advanced LLMs like GPT-4 and Gemini demonstrate remarkable reasoning, comprehension, and problem-solving capabilities, enabling applications in specialized domains such as law [2], medicine [3, 4], and education [5].
However, their closed-source nature raises concerns about data privacy, customization, and accessibility, especially for organizations seeking domain-specific solutions [6, 7]. Open-source models like LLaMA-3 [8] and Phi-4 [9] provide flexible alternatives, offering fine-tuning and in-context learning capabilities that often rival proprietary systems and exceed baseline results [10, 11].
A key candidate application of LLMs is LJP, which predicts judicial outcomes based on factual case details [12]. Currently, legal experts rely on extensive training to identify relevant laws, determine charges, and issue judgments [13]. Legal decision-making is inherently complex, requiring analysis of vast case-specific information while ensuring consistency and fairness [14]. This process remains labor-intensive, time-consuming, and prone to bias [14].
@@
To address these challenges, this work develops a tailored solution for Arabic LJP. Specifically, we:
@@
Collect and preprocess Arabic legal case data from the Saudi commercial court222https://laws.moj.gov.sa/ar/, creating a domain-specific dataset.
@@
Prepare a fine-tuning dataset by crafting 75 diverse Arabic instructions of varying lengths, complexity, and styles to enhance model generalizability and simulate human-centric prompting scenarios.
@@
Benchmark open-source LLMs, including LLaMA-3.2-3B-instruct and LLaMA-3.1-8B-instruct, under zero-shot, one-shot, and fine-tuning settings using LoRA.
@@ II Related Work
In recent years, Legal Judgment Prediction has attracted substantial attention, driven by advancements in NLP techniques [18]. Early approaches to LJP primarily utilized rule-based systems [19], relying on predefined legal rules and templates to predict outcomes. These systems were later replaced by traditional ML models with manually crafted features, introducing statistical methods that improved performance [20]. The emergence of neural-based approaches [21], particularly recurrent neural networks (RNNs) [22] and transformer-based architectures [23], revolutionized the field by enabling models to better capture the semantic complexity of legal texts.
Transformer-based models have proven transformative for legal text analysis by modeling intricate semantic and syntactic relationships. The BERT model [24], along with its specialized versions such as Legal-BERT [25], Legal-RoBERTa [26], CaseLawBERT, and InCaseLawBERT [27], has demonstrated exceptional performance in extracting semantic richness from legal documents. For instance, [28] showcased BERT’s effectiveness in classifying European Court of Human Rights cases, while [29] fine-tuned BERT models on the VerdictVaultPK dataset of Pakistani legal case judgments, achieving over 20 F1 points higher than traditional classifiers like TF-IDF and logistic regression. Similarly, [30] applied BERT to predict decisions in Indian legal judgments [31] and suggested the use of LLMs to address BERT’s context-length limitations.
Moreover, researchers have explored domain-specific adaptations of transformers for LJP in other languages. [32] demonstrated the importance of specialized transformer models for handling long-form Romanian legal judgments, emphasizing that general-purpose models often lack the contextual depth and domain knowledge required for legal tasks. In criminal law, [33] applied optimized BERT variants, outperforming sparse text representation models in predicting criminal case outcomes. Despite these advancements, challenges in multilingual and low-resource legal NLP persist. [34] noted that LLMs excel in high-resource languages like English but struggle with low-resource, non-Latin languages, a finding corroborated by [35].
@@
Interestingly, English-written instructions often perform strongly even for non-English tasks. For example, [17] found that non-native English instructions outperformed native Arabic ones in certain Arabic NLP tasks, demonstrating the adaptability of LLMs when leveraging high-quality, generalized instruction data in low-resource settings. Furthermore, [36] demonstrated that Llama3’s performance improved significantly when fine-tuned on high-quality Arabic datasets specifically curated for social media tasks.
Additionally, [37] utilized LLaMA, GPT, and Jais for Arabic LJP, showing that translating legal cases into English significantly improved model performance. Their findings further emphasize the need for Arabic-centric models tailored to the legal domain.
Recently, the potential of LLMs for LJP has been further explored. [38] and [39] benchmarked GPT-3.5 and LLaMA-2 on Indian court datasets, while [40] evaluated GPT-3’s performance on the Uniform Bar Examination, showcasing its capability in legal reasoning and generating contextually relevant answers. Transformer models like LegalBERT and T5 have also excelled in structured text generation, classification, and summarization tasks [41, 42], further solidifying their role in automating and enhancing legal systems.
Despite significant advancements in legal NLP, Arabic LJP remains underexplored. Challenges such as the morphological complexity of Arabic, syntax variability, and the scarcity of high-quality annotated datasets have hindered progress. Existing Arabic corpora are often fragmented, domain-specific, or poorly annotated, limiting the development of robust models.
@@ III-A Dataset Collection and Preparation
In this research, we constructed a domain-specific dataset for Arabic LJP by collecting court judgments from the publicly accessible Saudi Ministry of Justice Judgment Publication Platform333https://laws.moj.gov.sa/ar/. This platform hosts a substantial repository of legal cases, particularly within the domain of commercial law. Fig. 2 presents a sample data point.
@@
The embeddings were generated using a BERT-based sentence embedding model from Sentence Transformers [43], specifically the paraphrase-multilingual-MiniLM-L12-v2 model444https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2. This model supports multiple languages, including Arabic, and enables efficient semantic retrieval of relevant legal judgments.
This following table illustrates how the model processes legal instructions, facts, and reasons to predict judicial outcomes. It highlights the importance of clear prompts and structured data in enhancing LLM performance for Arabic LJP.
III-B Fine-Tuning
We use LoRA (Low-Rank Adaptation) [44], a parameter-efficient fine-tuning method enabling large pre-trained models to adapt with minimal computational cost. LoRA lowers memory usage while maintaining performance, allowing fine-tuning on resource-limited hardware for large-scale tasks.
Both models were trained for two epochs on a single A100 GPU, using a batch size of 32 and a maximum sequence length of 2048. LoRA was applied with a rank of , a scaling factor of , and a dropout rate of 0.1. The target modules included q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, and down_proj, all with a dropout of 0.1. Training was performed for one epoch using the AdamW optimizer [45] with a learning rate of .
@@ III-C Evaluation Metrics
For quantitative evaluation, we employed ROUGE-F [46], BLEU [47], and BERTScore [48] which are standard metrics for text generation tasks:
@@
BERT-score-F: Computes the F1-score using contextual embeddings to measure semantic similarity, making it well-suited for LJP, where precise legal understanding is crucial.
@@ III-E Model Inference
The models were deployed for inference using vLLM [51] on four NVIDIA A16 GPUs. The maximum sequence length was set to 2048 tokens. The temperature parameter was fixed at zero to ensure reproducibility and deterministic outputs.
@@ IV Experiments and Results
IV-A LLM-Based Evaluation Analysis
Fine-tuning led to a substantial improvement in objective performance across all evaluated metrics. For the smaller Llama-3.2-3B model, fine-tuning resulted in an average score increase from 3.01 to 7.13, yielding a 4.14-point improvement across all metrics. In comparison, under in-context learning settings, the 1-shot performance of Llama-3.2-3B outperformed the baseline by an average of 1.65 points.
For the larger Llama-3.1-8B model, fine-tuning yielded a 1.55-point improvement, surpassing both its untrained counterpart and the fine-tuned Llama-3.2-3B. Across all qualitative evaluation criteria—including Coherence, Brevity, Legal Language, Faithfulness, Clarity, and Consistency—both in-context learning and fine-tuning consistently outperformed the untrained models, demonstrating the effectiveness of these adaptation strategies.
IV-B Standard Evaluation Analysis
The base models exhibited poor performance when evaluated on standard metrics. For instance, Llama-3.2-3B achieved ROUGE-1: 0.08, ROUGE-2: 0.02, BLEU: 0.01, and BERTScore: 0.54. The larger Llama-3.1-8B performed slightly better, with ROUGE-1: 0.12, ROUGE-2: 0.04, BLEU: 0.02, and BERTScore: 0.58, highlighting the inherent complexity of the task.
1-shot learning significantly improved performance, yielding an absolute increase of 6% in BERTScore (0.64), 9% in BLEU, and 15% in ROUGE-2. Our fine-tuned models demonstrated comparable performance, with an average difference of less than 3%, favoring the larger Llama-3.1-8B-FT. Our best-performing model achieved a BERTScore of 0.76, indicating strong similarity and highlighting the model’s robustness. Similarly, it attained a BLEU score of 0.26, ROUGE-2 of 0.41, and ROUGE-1 of 0.53.
| Metric | LLaMA-3.2-3B | LLaMA-3.1-8B | LLaMA-3.2-3B-1S | LLaMA-3.2-3B-FT | LLaMA-3.1-8B-FT |
| Coherence | 2.69 | 5.49 | 4.52 | 6.60 | 6.94 |
| Brevity | 1.99 | 4.30 | 3.76 | 5.87 | 6.27 |
| Legal Language | 3.66 | 6.69 | 5.18 | 7.48 | 7.73 |
| Faithfulness | 3.00 | 5.99 | 4.00 | 6.08 | 6.42 |
| Clarity | 2.90 | 5.79 | 4.99 | 7.90 | 8.17 |
| Consistency | 3.04 | 5.93 | 5.14 | 8.47 | 8.65 |
| Avg. Qualitative Score | 3.01 | 5.89 | 4.66 | 7.13 | 7.44 |
| ROUGE-1 | 0.08 | 0.12 | 0.29 | 0.50 | 0.53 |
| ROUGE-2 | 0.02 | 0.04 | 0.19 | 0.39 | 0.41 |
| BLEU | 0.01 | 0.02 | 0.11 | 0.24 | 0.26 |
| BERT | 0.54 | 0.58 | 0.64 | 0.74 | 0.76 |
@@ V Discussion
This section examines the performance of LLM configurations in LJP, focusing on the effects of instruction generalization, fine-tuning, and in-context learning. We analyze model behaviors, highlight strengths and limitations, and discuss trade-offs in adapting LLMs for this specialized task.
V-A Instruction Generalization Analysis
To evaluate the impact of fine-tuning on model performance and its ability to generalize across a diverse range of instructions, we compared two sets of models: Llama-3.2-3b-instruct and Llama-3.2-3b-finetuned, as well as Llama-3.1-8b-instruct and Llama-3.1-8b-finetuned. The models were trained on a set of 75 diverse instructions to encourage generalization and promote human-centric behavior.
We measured model performance using the BERT F1 score and applied the Wilcoxon signed-rank test to assess statistical significance, with an alpha level set to 0.05. The significance test was applied as follows: the 75 instructions were distributed across the test set, and for each instruction, we computed the BERT F1 score. The individual scores were then aggregated by taking the average BERT score for each model, before and after fine-tuning. The Wilcoxon signed-rank test was then performed on the 75 averaged BERT scores to determine whether the improvements in performance post-fine-tuning were statistically significant.
The results demonstrated significant improvements in performance after fine-tuning. Specifically, the Llama-3.2-3b-instruct model showed a 20% absolute increase, from 0.54 to 0.64, with a p-value of 0, indicating a statistically significant enhancement. Similarly, the Llama-3.1-8b-instruct model exhibited an 18% absolute improvement, from 0.58 to 0.74, with a p-value of 0, confirming statistical significance.
These findings underscore the effectiveness of training on a diverse set of instructions to improve instruction generalization and achieve more human-centric performance. The observed improvements in both models after fine-tuning demonstrate the models’ enhanced ability to generalize across instructions, while ensuring they remain aligned with human-centric goals.
V-B Finetuning vs. In-Context Learning
As shown in Table I, finetning using LoRA significantly outperformed in-context learning across all metrics, Furthermore, in-context learning requires the model to manage extended context sizes, which becomes increasingly costly in legal domain due to the inherently large and complex nature of legal cases. Moreover, the necessity to embed and store extensive training examples further escalates the computational and storage requirements over time, posing a scalability challenge for practical deployment in real-world legal applications.
Given these considerations, finetuning emerges as the more pragmatic approach. It not only delivers significantly enhanced performance but also maintains a moderate context size, thereby reducing computational overhead and facilitating easier deployment. These findings suggest that a fine-tuning strategy, especially using efficient methods like LoRA, is better suited for advancing legal NLP applications.